OpenStudy (anonymous):

The four sides of a garden are 14 1/2 feet, 31 2/3 feet, 28 1/3 feet, and 22 1/2 feet. How many feet of fencing are needed to go around the garden? First estimate the answer. Then solve for the exact answer.

OpenStudy (ddcamp):

For estimating, I would round to the nearest foot: 14 1/2 → 14 31 2/3 → 32 28 1/3 → 28 22 1/2 → 23 Estimate sum = 14+32+28+23 = 97 feet

OpenStudy (ddcamp):

For the exact answer, I would start by adding up the whole number parts: 14 + 31 + 28 + 22 = 95 Then find the LCD on the fraction parts (= 2×3 = 6) and convert to common denominators and add: 3/6 + 4/6 + 2/6 + 3/6 = 12/6 = 2 And add the whole number part and the fraction part: 95 + 2 = 97 feet.

OpenStudy (anonymous):

so i put 97 for the estimated and 97 for the exact answer

OpenStudy (ddcamp):

Yes. There are different ways to estimate though (rounding to the nearest 5 or 10), so you could get a different estimation.

OpenStudy (anonymous):

thanks can u help me with another one